# Assurance of AI Systems From a Dependability Perspective

### v2

@@ Abstract
A key element in this “dependability” perspective is a requirement to have near-complete understanding of the behavior of critical components, and this is considered infeasible for AI and ML. Hence the dependability perspective aims to minimize trust in AI and ML elements by using “defense in depth” with a hierarchy of less complex systems, some of which may be highly assured conventionally engineered components, to “guard” them. This may be contrasted with the “trustworthiness” perspective that seeks to apply assurance to the AI and ML elements themselves.
@@
For guards that require perception using AI and ML, we examine ways to minimize the trust placed in these elements; they include diversity, defense in depth, explanations, and micro-ODDs. We also examine methods to enforce acceptable behavior, given a model of the world. These include classical cyber-physical calculations and envelopes, and normative rules based on overarching principles, constitutions, ethics, or reputation.
We apply our perspective to autonomous systems, AI systems for specific functions, general-purpose AI such as Large Language Models, and to Artificial General Intelligence (AGI), and we propose current best practice and an agenda for research.
1 Introduction: Assurance for Traditional Systems
@@
“If a builder build a house for some one, and does not construct it properly, and the house which he built fall in and kill its owner, then that builder shall be put to death” [69, Section 229].
@@
Systems with control and procedural mechanisms implemented by computers (so-called cyber-physical systems, CPS) drive the state of the art in safety engineering and assurance today [102]. The central tenet of the approach taken, which we call the dependability perspective, is that there must be near-complete understanding of how the given system works, what are its hazards, how these are eliminated or mitigated, and how we can be sure all this is done correctly. The evidence and arguments that justify confidence in the claims documenting this understanding constitute what is called an assurance case (see later for details and references).
@@
Systems that use Artificial Intelligence (AI) and Machine Learning (ML) are both an evolution and a step change from their predecessors: although they often automate existing systems and procedures, they work in different ways to what has gone before, and they can also do different things. Because they work in different ways, it is difficult to apply established methods for safety engineering and assurance to AI and ML, even when they are used in familiar or slightly extended contexts, such as automated control and autonomous systems. In particular, the behavior of systems based on ML is developed experimentally, so their inner workings are opaque and do not support the understanding required for justified confidence.
And because they can be used to do new things, some AI and ML systems introduce entirely new hazards: potentially, they can substitute for people in activities that previously required human levels of perception, language, intelligence, and judgement, and where failure can go beyond physical harm and can affect personal wellbeing, relationships, and society at large. Beyond this fairly incremental progression lies the step to Artificial General Intelligence (AGI), with potentially superhuman performance on significant activities, plus imagination, agency with independent goals, and possibly consciousness.
Our aim in this report is to identify and briefly describe issues, possible methods, and difficulties in assurance for systems with significant AI and ML content. We do this mainly from the dependability perspective, where AI and ML are accommodated by guarded architectures that use “guards” or “monitors” to provide assurance for the overall system with few or no assumptions on the AI and ML components.
@@
The dependability perspective may be contrasted with what we call the trustworthiness perspective, which does claim some assurance in the behavior of AI and ML components that have been developed, analyzed, tested, augmented, or restricted in various ways. [^1] Terminology across different fields is always difficult and sometimes contentious; the field of dependability regards “trustworthiness” as a synonym for “dependability,” whereas AI tends to use “safety” as the required property (e.g., the UK Government has an AI Safety Institute) and “trustworthiness” as the means for achieving and assuring it. The adjective “trustworthy” carries a somewhat anthropomorphic tone that we discuss in Section 4 and so we generally prefer the more neutral, engineering terminology of dependability. However, we need a name for the alternative perspective and we use “trustworthy” for that purpose. Both perspectives have merit and in practice there is a continuum or spectrum between them. In particular, we envisage guarded architectures that are recursively structured where “first level” guards might use some AI and ML and themselves be guarded by simpler systems, eventually bottoming out on conventionally engineered and assured guards so that the overall architecture provides “defense in depth.” The architecture and its assurance will vary according to how much assurance “credit” is taken for trustworthiness or diversity of AI and ML components [26]. We call this the dependability/trustworthiness spectrum; a “pure” dependability perspective takes no credit for trustworthiness of AI and ML components, while a “pure” trustworthiness perspective claims full assurance credit for those components.
@@ 1.1 Traditional Systems and their Assurance
State of the art non-AI cyber-physical systems such as aircraft flight control, safety systems such as nuclear shutdown, and all manner of systems within critical infrastructure, medical devices, personal gadgets and much else are generally engineered and assured for suitably high levels of safety and other required attributes, such as security or effectiveness, all generically referred to as dependability [105, 81]. In outline, the process for doing this begins with identification of the potential hazards that the proposed system might entail. A hazard is a circumstance with an unacceptably high risk of leading to harm or other undesired outcome. Hazard analysis is conducted in the context of assumptions about the environment in which the system will operate and is not an exact science: even its most effective methods can be imperfect and their application requires skill, knowledge and experience [119, 80].
@@
After some rounds of iteration on hazard identification and modifications to the system goals and design, we will have a set of requirements for the critical desired behavior of the computer control system that, with high confidence, ensures dependability of the overall system in its assumed environment. Identification and articulation of properties assumed about the environment are fundamental to formulation of requirements and are often the most difficult and fault-prone aspects of the entire system engineering endeavor. The analysis and reasoning that shows that the requirements ensure safety and other critical properties within their environment is an assurance task that we term dependability requirements validation.
@@
We then implement the system according to its specifications and architecture. The mechanisms that ensure an architecture is faithfully represented in the system implementation are among the most difficult engineering challenges in computer science (involving operating systems, “buses,” distributed consensus, state-machine replication, transaction mechanisms etc.) and should employ only well-attested techniques and products with no “homespun” solutions [145]. During implementation, we may discover new hazards and the whole process iterates: the new hazards cause revision to the requirements [^2] Confusingly, these revisions are often called derived requirements (the term comes from avionics); it is confusing because their essence is that these requirements were not derived during the main process of requirements development. and their safety validation, and also to the specifications and hence to the implementation.
@@
Any part of the implementation that is not derived from the requirements must be shown to have no unacceptable impact. [^3] Software libraries provide an example of implementation content that is not derived from requirements: we might require only the trigonometric functions, but the whole library is installed as part of the implementation.
@@
Each of the assurance validation and verification tasks states that some properties “must be shown” to hold; by this, we mean that there must be reasons why the properties hold, and these reasons must be clearly articulated and justified. The state of the art for doing this is an assurance case (a generalization of safety cases [91, 1]) that provides an organized presentation based on claims, evidence, and argument [24, 147]. Claims identify properties of the system and/or its environment; evidence refers to observations, measurements, or experiments on the system or its means of construction or on its environment that justify certain claims; and the argument uses the evidence to establish a hierarchy of claims culminating in a significant top claim. The arguments of an assurance case are not free form but structured as hierarchy of argument steps, each of which establishes a “parent” claim on the basis of one or more “child” claims (we usually say subclaims) established at lower levels, or by evidence. A portion of a graphical rendering of an assurance argument is displayed in Figure 1; our preferred treatment of modern assurance cases, which we call Assurance 2.0, is presented elsewhere [28] and builds on the ideas that a strong assurance case should be indefeasible [148, 29], based on established theories [171], and subjected to dialectical examination [27].

### v3

@@ Abstract
On its own, testing is insufficient for assurance when very high levels of confidence are required. Hence, a key element in the “dependability” perspective is a requirement to have thorough understanding of the internal design and operation (and hence behavior) of critical system components and their interaction. This is considered infeasible for AI and ML because their internal operation is developed experimentally over a limited (albeit large) set of training examples and is opaque to detailed understanding. Hence the dependability perspective, as we apply it here, aims to minimize trust in AI and ML elements by using “defense in depth” with a hierarchy of less complex systems, some of which may be highly assured conventionally engineered components, to “guard” them. This may be contrasted with what we call the “trustworthiness” perspective that seeks to apply assurance to the AI and ML elements themselves by various forms of careful training, fine tuning, internal “guardrails” and automated examination.
@@
For guards that require perception using AI and ML, we examine ways to minimize the trust placed in these elements; they include diversity, defense in depth, explanations, and micro-ODDs (Operational Design Domains). We also examine methods to enforce acceptable behavior, given a model of the world. These include classical cyber-physical calculations and envelopes, and normative rules based on overarching principles, constitutions, ethics, and reputation.
We apply our perspective to autonomous systems, AI systems for specific functions, general-purpose AI such as Large Language Models (LLMs), and Artificial General Intelligence (AGI), and we propose current best practice and conclude with a fourfold agenda for research in which we recommend development and application of: a) new methods for hazard analysis suited to AI systems; b) layered recursively structured architectures for runtime verification and defense in depth; c) assurance for AI-based perception, and d) improved understanding of human and machine cognition, shared intentionality, and emergent behavior.
1 Introduction
Much discussion of potential risks with AI concerns “existential” threats [101], but we suggest that lesser—yet significant and widespread—hazards will arise as near-term AI is embedded in other systems. Hence, it is an urgent task to provide users of systems incorporating AI, and also those responsible for their safety and security, with methods to assess their hazards and to incorporate appropriate means of mitigation, such as architectures that provide assured runtime checking with diversity and defense in depth. In addition, AI developers need to be aware of these issues and should provide suitable APIs and other mechanisms that enable integration of AI within assured systems.
By “assured” we mean that there are good reasons to believe that a system will not exhibit certain harmful behaviors or will do so very rarely, where “rarely” will be quantified as, say, no more than once in a million demands. “Assurance” is the process of justifying such claims, so that it is rational to believe them and highly likely that they will be borne out by subsequent experience. Assurance is normally required only for systems whose harms are such that they must be rare. Thus assurance is associated with methods for building high quality or dependable systems and discussions of assured systems generally (as here) have two interrelated threads: a) how to build systems that are dependable, and b) how to justify confidence that they are so.
Risk owners in critical industries are responsible for understanding, assessing and ensuring that hazards (i.e., circumstances with unacceptable risk of leading to harm) are addressed according to existing national and international frameworks. Therefore, they need methods for analyzing the potential harm (and benefit) from AI and they need to develop justifications (e.g., as assurance cases [39]) for the safety and security of systems that include AI—and their regulators will need to understand and assess these. We also note that unregulated industries still have overarching responsibilities to address hazards and vulnerabilities. Recently, many countries have established government agencies such as the UK AISI (aisi.gov.uk) to support risk owners by developing assessment methods and also by providing insights on AI systems that may be components in these critical applications.
There is much recent work to improve the reliability of AI-based systems, and we welcome these developments. However, although much of this work has been successful, it does not amount to assurance, which requires a rationally justified claim that failures will not exceed some bound. (Risk is generally understood as the product of severity of a harm and its frequency. If severity is assumed to be fixed, we limit the risk by establishing a bound on its frequency, that is on failures with respect to the assured claim.) The contribution of this report is to discuss the assurance challenges of AI and to explain and explore classic strategies for achieving trustworthy dependable systems, where dependable is an umbrella term encompassing safety, security, reliability, and so on.
The field of dependable systems engineering employs a standard terminology [150]. Specifically, failure occurs when a system departs from its explicitly specified or implicitly expected behavior. The cause of a failure is a fault, which may be the failure of some subsystem or component, or an oversight or mistake in the design or implementation of the system. Faults precipitate errors, which are incorrect values or configurations of the system’s internal state. It may be possible to detect and correct errors before they propagate to cause failure, thereby providing fault tolerance. And failure need not always be terminal: there may be some external processes for recovery or adaptation that provide resilience and allow (some aspects of) the larger context to continue.
Figure 1: Four State Model for Dependability
A 4-state model shown in Figure 1 portrays transitions on this ontology, corresponding to different circumstances and strategies for dependable operation [1, 43]. These transitions illustrate the three classic strategies for trustworthy dependable systems.
Fault avoidance.
This aims to eliminate transitions from the OK to Error state and is the domain of highly rigorous software engineering.
Fault tolerance.
Faults correspond to transitions from the OK to Error state, and Fault Tolerance to the transition back to OK.
Failure Management and Resilience.
Unsuccessful or missing fault tolerance allows transitions from the Error to Failure states, which we divide into Safe and Unsafe according to their consequences. Failure Management tries to ensure the transitions are to Safe Failures (e.g., a failed train comes to a halt) and provides some recovery process that enables transition back to the OK state (e.g., the train is taken out of service). There will be some loss associated with the failure and recovery.
Resilience corresponds to the transition from Unsafe Failure back to OK. It may be associated with considerable loss due to both the failure (e.g., a train crash) and the recovery or adaptation (e.g., revision to the signaling system).
These strategies are not mutually exclusive: dependable systems should employ a combination, with emphasis varying according to the information that is available, the nature of the technology being used, and the consequences of failure. For systems employing AI and ML we assess the three strategies as follows.
Fault avoidance.
As we describe in the following subsections, this requires strong understanding of the system context and requirements, and a means of assuring that these are applied and implemented correctly. Testing is an important means of assurance but on its own it is insufficient for anything more than everyday risks. We claim that the necessary understanding and assurance is currently infeasible for AI and ML and consequently fault avoidance cannot be the main dependability strategy for AI and ML systems. Proposals for “Guaranteed Safe AI” do exist [66], but we consider them speculative.
Fault tolerance.
This is a key strategy when components are relatively unreliable or hard to evaluate, but it does depend on having well-defined safety/security requirements (so that we can recognize transitions into the error state). It is the main strategy discussed in Sections 2 and 3.
Failure Management and Resilience.
This becomes an important strategy in circumstances where some failures are acceptable. It is particularly appropriate for new technologies or those used in ways not previously envisaged, where some degree of learning by doing is acceptable. Resilience becomes a major strategy in Sections 4 and 5.
However, it is not directly applicable to systems that are too important to fail, such as those where the risks are truly existential, irreversible, or otherwise of high consequence to society. Nonetheless, the technology of resilience could be applied to early warning and near-miss detection for these systems so that effective controls can be imposed before they reach a critical stage of development or deployment.
1.1 The Dependability Perspective on Assurance
@@
“If a builder build a house for some one, and does not construct it properly, and the house which he built fall in and kill its owner, then that builder shall be put to death” [97, Section 229].
@@
Systems with control and procedural mechanisms implemented by computers (so-called cyber-physical systems, CPS) drive the state of the art in safety engineering and assurance today [144]. The approach taken, which we call the dependability perspective, has fault avoidance as its primary strategy and fault tolerance, generally for hardware failures, as its secondary strategy, with resilience as a last resort. A consequence of this approach is that fault avoidance has to be assured to an extremely high level: typically less than one failure in a billion demands. Testing is insufficient for this (see Section 1.3) and must be supported by near-complete understanding of how the given system works, what are its hazards, how these are eliminated or mitigated, and how we can be sure all this is done and implemented correctly. The evidence and arguments that justify confidence in the claims documenting this understanding constitute what is called an assurance case (see later for details and references).
So far we have considered systems where component faults are the main cause of failures. However, failures in complex systems are often due to unanticipated interactions among components that are working correctly (i.e., according to their requirements, which may themselves by faulty). These are called system failures (as opposed to component failures) and many of the most egregious recent failures are of this type (or a combination, where a component fault escalates to catastrophe by precipitating a system failure). Examples include the nuclear accidents at Chernobyl, Three Mile Island, and Fukushima; airplane crashes such as Air France 447 and the 737 MCAS; spacecraft explosions such as Challenger and Ariane 5; marine disasters such as Deep Water Horizon and the Norwegian frigate “Helge Ingstad”; financial failures such as Long Term Capital Management (LTCM) and the banking collapse of 2008; miscarriages of justice such as the UK Horizon scandal, and software outages such as SolarWinds and CrowdStrike. System failures were famously identified by Perrow as “Normal Accidents” [185]. A related notion, due to Per Bak [15], is “Self-Organizing Criticality” (SOC), which can be seen as a “hidden hand” behind Normal Accidents [157].
System failures often result from drawing the system boundary too narrowly: the system is considered to comprise the “mechanism” and its immediate environment, including those who directly interact with it, but the wider socio-technical context [18] is overlooked. Modern system safety engineering and methods of hazard analysis such as Leveson’s System Theoretic Process Analysis (STPA) aim to identify and overcome these wider sources of system failure [153, 154]. An important consequence of AI is that it provides capabilities (such as natural language) that cause it to become much more deeply embedded in its socio-technical context than may be anticipated or recognized and thereby extend the system boundary. For example, Jatho and colleagues applied STPA to an ML-driven prescription drug monitoring system and to a face recognition system used in criminal justice and found previously unrecognized socio-technical hazards in both cases [109].
We directed attention to the drawing of system boundaries when using AI and ML in a recent paper [43] and we will return to this topic in Section 5. However, current AI and ML systems are sufficiently unreliable that they are a significant source of component failure in systems that employ them. Consequently, we will focus on failures of AI and ML components, but the larger context should be kept in mind.
@@
Systems that use Artificial Intelligence (AI) and Machine Learning (ML) are both an evolution and a step change from their predecessors: although they often automate existing systems and procedures, they work in different ways than what has gone before, and they can also do intriguing new things. Because they work in different ways, it is difficult to apply established methods for safety engineering and assurance to AI and ML, even when they are used in familiar or slightly extended contexts, such as automated control and autonomous systems. In particular, the behavior of systems based on ML is developed experimentally, so their inner workings are opaque and do not support the understanding required for justified confidence. And because they can do new things, and appear to do them well enough to substitute for humans in some circumstances, AI and ML are being used in applications where they introduce entirely new hazards: by substituting for people in activities that previously required human levels of perception, language, intelligence, and judgment, failure can go beyond physical harm and can affect personal wellbeing, relationships, and society at large.
Beyond these fairly incremental progressions lies the step to Artificial General Intelligence (AGI), with potentially superhuman performance on significant activities, plus imagination, agency with independent goals, and possibly consciousness.
Our aim in this report is to identify and briefly describe issues, possible methods, and difficulties in assurance for systems with significant AI and ML content. We do this mainly from the dependability perspective: because we do not believe that AI and ML can provide well-assured fault avoidance, we accommodate them by fault-tolerance strategies using guards and/or diverse replicas or backups that monitor their behavior at runtime, often within a larger architecture that provides defense in depth and resilience. These architectures can deliver strong assurance for the overall system with only weak assumptions on the behavior of AI and ML components. Furthermore, the concept can be replied recursively within a system, so the AI and ML components can themselves have internal “guard rails,” and also externally to the wider system and organizational environment.
We are not the first to advocate these methods. Notably, the interim report from the Seoul Summit on Safety of Advanced AI identifies a broad range of risks and issues and highlights the “Swiss Cheese Model” of protection [21, Section 5.1.2], which can be seen as a combination of diversity and defense in depth.
@@
The dependability perspective may be contrasted with what we call the trustworthiness perspective, which does claim assurance (i.e., fault avoidance) for the behavior of AI and ML components that have been developed, analyzed, tested, augmented, or restricted in various ways [234]. [^1] Terminology across different fields is always difficult and sometimes contentious; the field of dependability regards “trustworthiness” as a synonym for “dependability” [10] and “safety” as a particular case within dependability, whereas AI tends to use “safety” as the generally required property and “trustworthiness” as the means for achieving and assuring it. The adjective “trustworthy” carries a somewhat anthropomorphic tone that we discuss in Section 4 and so we prefer the more neutral, engineering terminology of dependability. But then we need a name for the alternative perspective and we use “trustworthy” for that purpose. One criticism of the trustworthiness perspective is that it often fails to assess potential harms realistically and the confidence needed in their elimination. Nonetheless, both perspectives have merit when well executed and in practice there is a continuum or spectrum between them. In particular, we envisage guarded architectures that are recursively structured where “first level” guards might use some AI and ML (e.g., for perception) and themselves be guarded by simpler or diverse systems, eventually bottoming out on conventionally engineered and assured guards so that the overall architecture provides defense in depth.
The architecture and its assurance will vary according to how much assurance “credit” is taken for trustworthiness of AI and ML components [37]. We call this the dependability/trustworthiness spectrum; a “pure” dependability perspective takes no credit for trustworthiness of AI and ML components, while a “pure” trustworthiness perspective claims full assurance credit for those components. Historically, the dependability perspective, and its methods, are very similar to those developed several years ago for using “Commercial Off The Shelf” (COTS) and “Software Of Uncertain Pedigree” (SOUP) components within critical applications (e.g., non-ASIL software components in cars) [32, 190].
@@ 1.2 Traditional Systems and their Assurance
State of the art non-AI cyber-physical systems such as aircraft flight control, safety systems such as nuclear shutdown, and all manner of systems within critical infrastructure, medical devices, personal gadgets and much else are generally engineered and assured for suitably high levels of safety and other required attributes, such as security or effectiveness, all generically referred to as dependability [114]. [^2] Strictly, security is distinguished from dependability [10]: the former corresponds to the impact of the environment on the system, whereas the latter is the impact of the system on the environment. For our purposes, we can lump them together. In outline (a slightly more extended account is provided at [43]), the process for doing this begins with identification of the potential hazards that the proposed system might entail. A hazard is a circumstance with an unacceptably high risk of leading to harm or other undesired outcome (the corresponding concept in security is threat). Hazard analysis must consider more than the consequences of simple component failures, it must consider malfunction and unintended function, and also unexpected interactions among elements that are performing as intended (recall the earlier discussion of system failures). Methods of hazard analysis often build on previous experience and may need to be extended for new technology such as AI. For example, HAZOP [171] uses guidewords and asks “what might happen if this output is late/wrong/absent” and so on. AI may introduce new kinds of error so that the guidewords may need to include phrases such as “is a lie,” “is biased” or “is offensive.”
With systems that do new things, or that operate in challenging environments, there may be little relevant experience to guide hazard analysis, so it is often supported by experiments (e.g., prototyping or simulation). In cars, for example, this is termed “Safety of the Intended Functionality” (SOTIF) [113]; critics suggest augmenting these (often massive, but still incomplete) experiments with formal methods as these can, in principle, examine all cases within some context [212].
Note that humans may be part of the system (e.g., as operators) and their fallibilities and vulnerabilities must be taken into account. Hazard analysis is conducted in the context of assumptions about the environment in which the system will operate (which again may include humans) and must consider (previously) unanticipated circumstances within this context and also the suitability of the assumptions. These are demanding tasks, and hazard analysis is not an exact science: even its most effective methods can be imperfect and their application requires skill, knowledge and experience [112].
@@
After some rounds of iteration on hazard identification and modifications to the system goals and design, we will have a set of requirements for the critical desired behavior of the computer control system that, with high confidence, ensures dependability of the overall system and accurately characterizes its assumed environment. Identification and articulation of properties assumed about the environment are fundamental to formulation of requirements and are often the most difficult and fault-prone aspects of the entire system engineering endeavor. The analysis and reasoning that shows that the requirements ensure safety and other critical properties within their environment is an assurance task that we term dependability requirements validation.
@@
We then implement the system according to its specifications and architecture. The mechanisms that ensure an architecture is faithfully represented in the system implementation are among the most difficult engineering challenges in computer science (involving operating systems, “buses,” distributed consensus, state-machine replication, transaction mechanisms etc.) and should employ only well-attested techniques and products, with no “homespun” solutions [204]. During implementation, we may discover new hazards and the whole process iterates: the new hazards cause revision to the requirements [^3] Confusingly, these revisions are often called derived requirements (the term comes from avionics); it is confusing because their essence is that these requirements were not derived during the main process of requirements development. and their safety validation, and also to the specifications and hence to the implementation.
@@
Any part of the implementation that is not derived from the requirements must be shown to have no unacceptable impact. [^4] Software libraries provide an example of implementation content that is not derived from requirements: we might require only some trigonometric functions, but the whole library is installed as part of the implementation.
@@
Each of the assurance validation and verification tasks states that some properties “must be shown” to hold; by this, we mean that there must be reasons why the properties hold, and these reasons must be clearly articulated and justified. The state of the art for doing this is an assurance case (a generalization of safety cases [128, 1]) that provides an organized presentation based on claims, evidence, and argument [35, 207]. Claims identify properties of the system and/or its environment; evidence refers to observations, measurements, or experiments on the system or its means of construction or on its environment that justify certain claims; and the argument uses the evidence to establish a hierarchy of claims culminating in a significant top claim. The arguments of an assurance case are not free form but structured as hierarchy of argument steps, each of which establishes a “parent” claim on the basis of one or more “child” claims (we usually say subclaims) established at lower levels, or by evidence. A portion of a graphical rendering of an assurance argument is displayed in Figure 2; our preferred treatment of modern assurance cases, which we call Assurance 2.0, is presented elsewhere [39, 44] and builds on the ideas that a strong assurance case should be indefeasible [208, 40], based on established theories [237], and subjected to dialectical examination [38]. Barrett and colleagues describe application of Assurance 2.0 to an AI safety case [17].